Given the constraints
A+B+C<or=24,B+C>or=8and A>or=0,C>or=0.
Maximize24-A-B

A:amount of time spent on schoolwork
B:amount of time spent on fun
C: amount of time spent on pay work
1
Expert's answer
2020-04-28T18:54:19-0400

here in the question only A and C are give as "A,C\ge0" . But the B is amount of time spent on fun and time also cannot be negative and it should be "B\ge 0."


This problem is converted to solve using simplex(Big-M) method and it is follows.

"maximize \ 24-A-B+0S_1+0S_2-Ma"

"subject\ to,\\
A+B+C+S_1\hspace{4 em}=24\\
B+C\hspace{4 em}-S_2+a\ =8\\
A,B,C\ge0"


iteration-1



a goes out from the basis and C comes in.

iteration-2

"since\ Z-C_j\ge 0"  optimum solution is occurred.

maximum is occurred when,

A=0,B=0,C=8 and Maximum is 24